Zamboni has broken down and is being pulled across the ice(consider a frictionless) The workers pull the Zamboni with a rope which makes an angle of 10.0with The horizontal force. The force of tension in the rope due to the pole of the two workers is 47.9 and if the repair bay is 40.0 M away will the workers be able to make it if they expended 2.00 x103 J of work?


1
Expert's answer
2021-05-26T13:00:09-0400

Mass of zamboni

"2Tcos\theta=mg"

"M=\frac{2Tcos\theta}{g}=\frac{2\times 47.9cos10}{9.8}"

"=9.627kg"

Required time

"W=2\times 10^3J"

"W=Fd"

"F=\frac{w}{d}=\frac{2\times10^3}{40}=50N"

"50=9.627\times\frac{40}{t^2}=t=\frac{\sqrt{9.627\times40}}{50}=2.875"

Time taken

The net force forward will be the sum of forward components of the two ropes

Each rope , "Fx=Fcos\theta=47.9cos10"

"=47.17N"

So net forward force, "F_{forward}=2\times 47.17N=94.34N"


"F=\frac{m\triangle V}{\triangle t}"

"94.34=m\times\frac{40.0}{t^2}=94.34=9.627\times\frac{40.0}{t}"

"t=\sqrt{\frac{9.627\times 40}{94.34}}=\sqrt{4.08seconds}=2.02 sec"

"\therefore" The workers will make it to the repair bay